MILLIMETER-WAVE SPECTROSCOPY OF METAL-CONTAINING MOLECULES: A DECADE OF BLOOD, SWEAT AND TELESCOPES

\underline{L. M. ZIURYS}, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

Metal-containing molecules are important in a wide variety of areas, including organic synthesis, catalysis, chemical vapor deposition and in biology. They also play a significant role in interstellar chemistry, where they trace the unusual physical and chemical processes during the later stages of stellar evolution. Therefore, high-resolution measurements of the gasphase spectra of metal-containing species impact on many fields, providing structural information, bonding characteristics and the necessary “rest-frequencies” for radio astronomical observations.

For over a decade, the Ziurys group has been recording pure-rotational spectra of transient metal-bearing species using millimeter/submillimeter direct absorption techniques. These molecules have been created through the gas-phase reaction of the metal vapor, produced by a modified Broida oven, with an appropriate precursor. Metals of particular interest have included lithium, sodium, magnesium, potassium, calcium, aluminum, and the 3d transition metals, combined with a variety of simple ligands such as OH, CN, CCH, CH_{3}, NH_{2}, and even carbon and nitrogen atoms. Several of the molecules investigated are common organic reagents such as LiCH_{3}, NaNH_{2} and CuCN; others are short lived free-radicals with high-spin ground states, e.g. MnF (^{7}Σ^{+}), CoCN (^{3}Φ_{i}) and NaC (^{4}Σ^{−}), some of which exhibit Renner-Teller splittings or quasi-linear effects in vibrationally excited states. A large fraction of these molecules are also of astrophysical interest and have been studied in interstellar/circumstellar gas. In this talk, the interesting chemistry and spectroscopy of these species will be discussed. In addition, implications of recent astronomical identifications will be summarized.